What is the pH of an aqueous solution of 0.006534 M Ba(OH)2 at 25oC? (Give your answer to two decimal places)


1
Expert's answer
2021-10-05T01:43:16-0400

[OH



]=0.006534×2=0.0131M




⇒pOH=−log(0.0131)=1.88




As we know that,




pH+pOH=14




∴pH=14−pOH=14−1.88=12.12